40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Jun 15 21:41:16 2012, frank, tara, Notes, DAQ, Temperature control updated 
    Reply  Thu Apr 4 18:45:38 2013, tara, Notes, DAQ, Temperature control updated 
Message ID: 1139     Entry time: Thu Apr 4 18:45:38 2013     In reply to: 988
Author: tara 
Type: Notes 
Category: DAQ 
Subject: Temperature control updated 

[peter, tara]The temperature servo for the chamber is back on, the current setup is at 31.2 C.

There was a problem with C3:PSL-VAC_CHAMBERTEMP channel, and I could not run the script for temperature control of the chamber. Peter helped me figure out what happened. It turned out that one of the parenthesis in the database file (cavities.db) was missing due to an accidental delete, and the name of the channel was too long (it was working before, I don't know why).

Anyway, the channel was renamed to C3:PSL-VAC_CHMBRTMP, in 1)cavities.db, 2)rcav_PID_2012_06_15.pl, and 3) medm screen for controlling the servo. The temperature servo is working again.


We switched the temperature readout channels used for temp feedback control to improve the signal. The new signal is significantly smoother.

      The signals from 4 thermostats around the vacuum chamber were acquired through 4 channels, C3:PSL-RCAV_SENSE(1-4). These channels were then connected to DAQ. This made the signal noisy because the resolution of analog to digital converter was low. In order to fix that we use an analog circuit to sum and average the signals from 4 sensors then amplify it before sending to DAQ,C3:PSL-RCAV_TEMP, then calibrated it to C3:PSL-VAC_CHAMBERTEMP by comparing RCAV_TEMP [V] to RCAV_TEMPAVG[C] which is calibrated to deg C already. 

      CHAMBERTEMP = (RCAV_TEMPx-0.495) + 34.957


     We corrected the perl script (in SUN machine) used for thermal feedback on the heater jacket. Now the script is named rcav_PID_2012_06_15.pl, see wiki. The servo is now back on.


ELOG V3.1.3-